Grade 12th passMechanics

A cube is placed so that one corner is at the origin and three edges are along the x-, y-, and z-axes of a coordinate system. Use vectors to compute (a) the angle between the edge along the z axes (line ab) and the diaganal from the origin to the opposite corner (line ad), and (b) the angle between line ac (the diaganal of a face) and line ad.

Anvith Reddy Nagulapally
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Let the cube have side = 1 

Then the coordinates for the corners are: 
a(0,0,0) , b(0,0,1), c(0,1,1) and d(1,1,1) 

The angle between two vectors u and v: u • v = IuI*IvI*cosθ or cosθ = (u • v)/ (IuI*IvI) 

a) ab • ad = • = 0*1+0*1+1*1 = 1 
....IabI = √(0²+0²+1²) = 1 , IadI = √(1²+1²+1²) = √3 
cosθ = 1/(1*√3) 
θ = 54.7° 

Now I'm sure you can work out b)...θ = 35.3°